OpenStudy (ttp):

Describe how the graph of y= x2 can be transformed to the graph of the given equation. y = (x - 15)2 Shift the graph of y = x2 right 15 units. Shift the graph of y = x2 left 15 units. Shift the graph of y = x2 down 15 units. Shift the graph of y = x2 up 15 units.

OpenStudy (brooke..help00):

I think the 3rd one

OpenStudy (mathmale):

Please write this squaring function as y=x^2; x2 is ambiguous. Similarly, write y=(x-15)^2. There are two main kinds of translation here: to the left or right, or up or down. y=(x-3)^2 has a graph which is that of y=x^2, except that it has been translated 3 units to the right. You need to know the effect of all four of the possibilities: y=(x+h)^2 y=(x-h)^2 y=x^2+h y=x^2-h Could you please look up "translations" in your algebra book, or your online textbook? Unfortunately, "the 3rd one" is not correct.